OpenStudy (anonymous):

how to solve if a=-15,b=+20 and c=-212 a3/b9+c67

OpenStudy (anonymous):

does a3 mean a^3 as in a "cubed" or "to the third power"?

OpenStudy (anonymous):

(-15)^3/(20)^9 + (-212)^67 = -7.319858 * 10^155

OpenStudy (anonymous):

(-15)*3/(20)*9 + (-212)*67 = -14224.25

OpenStudy (anonymous):

To solve, just substitute the letters for the numbers

OpenStudy (anonymous):

thanks